how would you invest $6400, part at 8% per annum and the remainder at 6% per annum so that at the end of one year the interest from the 8% invesment will be double the interest from the 6% invesment
Answer by lwsshak3(11628)

how would you invest $6400, part at 8% per annum and the remainder at 6% per annum so that at the end of one year the interest from the 8% invesment will be double the interest from the 6% invesment
..
let x=dollars invested at 8% per annum
then 6400-x = remaining dollars invested at 6% per annum
interest earned from 8% investment=twice that of 6% investment
.08x/.06(6400-x)=2
.08x=.12(6400-x)=768-.12x
.2x=768
x=768/.2=$3840
ans:
You would invest $3840 at 8% interest and $2560 at 6% interest.